OpenStudy (anonymous):

2- sqaure root of -27

OpenStudy (anonymous):

Are you asking to simplify \(2-\sqrt{-27}\)?

OpenStudy (anonymous):

write the complex number in standard form

OpenStudy (anonymous):

Well...\[2- \sqrt{-27}=2-\sqrt{-1}\sqrt{27}=\boxed{2-i\sqrt{27}}\]You could also write \(\sqrt{27}\) as \(3\sqrt{3}\) but that doesn't make it any more in standard form as what I wrote.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

thanks